OpenStudy (anonymous):

Which expressions are equal to 6 when x = 8 and y = 1? Choose all answers that are correct. A. x – y + 1 B. (x ÷ y) – 2 C. x + y – 5 D. xy – 2

OpenStudy (anonymous):

B is one of them

OpenStudy (anonymous):

plug in the values, and see if you get the answer 6

OpenStudy (anonymous):

:l I have for all of them, But I'm not sure about D

OpenStudy (anonymous):

We know that A is incorrect since 8−1+1=8, and not 6

OpenStudy (anonymous):

B is one of them

OpenStudy (anonymous):

We know B is correct since 8 divided 1 - 2 is 6

OpenStudy (anonymous):

we know c is incorrect since 8+1−5=4

OpenStudy (anonymous):

we also know that xy – 2 is correct since 8*1-2=6

OpenStudy (anonymous):

so B and D are correct

OpenStudy (anonymous):

Oh I did that last one wrong, Thanks for your help

OpenStudy (anonymous):

lol, no problem